OpenStudy (anonymous):

Find the sum of the first 6 terms of the geometric series whose first term is 81 and whose common ratio is −2/3

OpenStudy (anonymous):

48.5999?

OpenStudy (anonymous):

Series is as follows: 81,-54,36,-2416,-10 2/3 Total is 44 1/3 or 133/3 or 43.3333

OpenStudy (anonymous):

Thats -24,16

OpenStudy (anonymous):

how u get that

OpenStudy (anonymous):

series

OpenStudy (anonymous):

You get the first number which is 81 and apply the ratio to each subsequent number

OpenStudy (anonymous):

how do u apply it?

OpenStudy (anonymous):

Ratios are multipliers

OpenStudy (anonymous):

ok so multiply it by 81 then by that answer and so on tell i get to what?

OpenStudy (anonymous):

So 81*(-2/3) = -54 That's number 2 in the series. Repeat: 54*(-2/3) = 36 Etc.

OpenStudy (anonymous):

That's -54*(-2/3). Damn typos :)

OpenStudy (anonymous):

Then you tally the first 6 numbers in the series.

OpenStudy (anonymous):

ok i got it i think check my answers i will post

OpenStudy (anonymous):

Happy to see folks who want more than just an answer :)

OpenStudy (anonymous):

haha cool happy to see folks that will help and not just talk all day on the chat lol

OpenStudy (anonymous):

I'm supposed to be studying for my Chem and PolSci final. Distracting myself ;)

OpenStudy (anonymous):

ha well ill get back with what i get for the next but dont let me get in your way :)

OpenStudy (anonymous):

Geometric series:\[a+ar+ar^2...ar^n\]With a=81 and r=-2/3, The first six terms are:\[81+81(-2/3)+81(-2/3)^2...81(-2/3)^5\]This gives:\[81(1-2/3+4/9-8/27+16/81-32/243)\]Or,\[81(\frac{243-162+108-72+48-32}{243})=81\frac{133}{243}=\frac{10,773}{243}\]Or,\[\frac{3592}{81}\]

OpenStudy (anonymous):

Check my arithmetic tho :)

OpenStudy (anonymous):

ha

OpenStudy (anonymous):

3591 on top :)

OpenStudy (anonymous):

Eseidls expansion of the series is the best way to look at it. Depends on what math class you are in. For Calc 2 you'll want to be working through his process.

OpenStudy (anonymous):

i got the check answer one up

OpenStudy (anonymous):

agreed. I checked it using http://www.analyzemath.com/Calculators_4/geometric_series_calc.html